OpenStudy (anonymous):

Explanation? I know I need to take the derivative first just not sure how. Consider the function f(x) = xe^2x. Find a critical point. Does f have a relative maximum, a relative minimum, or neither at that point?

OpenStudy (emmatassone):

First take the derivative:\[f(x)=x.e ^{2x}\rightarrow f \prime(x)= e ^{2x}+x(2e ^{2x})= e ^{2x}.(1+2x)\] now we equal to zero the equation to find critical points:\[f \prime(x)=0 \rightarrow e ^{2x}(1+2x)=0\rightarrow 1+2x=0\rightarrow x=-\frac{ 1 }{ 2 }\] we used \[e ^{2x}\neq0 (\forall x \in \mathbb{R})\] finally evaluate x=-1/2 in the second derivative to see if its a maximum or minimum. Here its a graph to corroborate the solution.

OpenStudy (anonymous):

Can you explain how you got the derivative? Is there a formula or something to use?

OpenStudy (emmatassone):

there is a formula to derivative a multiplication:\[f(x)= g(x).h(x)\rightarrow f \prime(x) = g \prime(x).h(x) + g(x).h \prime(x)\] and there is also a formula to get the derivative of each funtion. you can look for them in internet

OpenStudy (anonymous):

Thanks!